# On solving basic equations over the semiring of functional digraphs

### v3

@@ Abstract
Endowing the set of functional digraphs (FGs) with the sum (disjoint union of digraphs) and product (standard direct product on digraphs) operations induces on FGs a structure of a commutative semiring $R$. The operations on $R$ can be naturally extended to the set of univariate polynomials $R[X]$ over $R$. This paper provides a polynomial time algorithm for deciding if equations of the type $AX=B$ have solutions when $A$ is just a single cycle and $B$ a set of cycles of identical size. We also prove a similar complexity result for some variants of the previous equation.
@@ 1 Introduction
A polynomial in $R[X_{1},X_{2},\ldots,X_{k}]$ represents a (infinite) set of functions which have a common substructure provided by the coefficients (modulo isomorphism). The factorization is an interesting inverse problem in this context. Indeed, assume to have a functional graph $G$. Can $G$ be factorized i.e., decomposed into the product or the sum (or a combination of the two) of functional graphs with smaller vertex set [^1] We stress that factorization here is meant in the context of abstract algebra which has nothing to do with the notion of factor graph often used in graph theory.? If this question has been largely investigated for general graphs, little is known for functional digraphs.
Let us consider another version of the factorization problem. Assume that we have partial information (or partial assumptions) on the factorization of $G$. This partial information is represented by a sequence of functional digraphs $A_{1},A_{2},\ldots,A_{k}$ and the question of factorization can be reformulated as ‘does the following equation
@@
admit a solution?’ And more generally one can ask for solutions of the following
@@
The interest of this result is manyfold. Without any doubt it will significantly improve the performances of the software pipeline in  [4] for the case in which one is interested to the existence of solutions and not to their enumeration. It will shed some light in some connected important questions concerning the cancellation problem for functional graphs i.e., the problem of establishing if $G\cdot F\approx G\cdot H$ implies $F\approx H$ where $F,G$ and $H$ are FGs and $\approx$ is graph isomorphism [5, 8]. We refer the reader to the classical book of Hammak for the factorization and the cancellation problems for more general graphs [6].
@@ 2 Background and basic facts
Figure 1: An example of polynomial equation over FGs (above) and its solution (below).
@@
It is therefore natural to ask if simpler equations admit solution algorithms with lower complexity.
@@ 3 Solving basic equations on permutations
From now on, we will focus on a subclass of functional graphs, namely the graphs of permutations. For this reason, we introduce a convenient notation called C-notation taken from [4]. We note $C^{1}_{p}$ the graph made by a single cycle of size $p$. Similarly, $C^{n}_{p}$ denotes a graph which is the sum of $n$ graphs $C^{1}_{p}$. According to this notation, it is clear that $C^{n_{1}}_{p}+C^{n_{2}}_{p}=C^{n_{1}+n_{2}}_{p}$ and $kC^{n}_{p}=C^{kn}_{p}$ for any natural $n_{1},n_{2}$ and $k$. With this notation, the classical result that the graph $G$ of a permutation is a union of disjoint cycles translates into
$G=\sum_{i=1}^{l}C_{p_{i}}^{n_{i}}$
for suitable positive integers $p_{1},\ldots,p_{l}$ and $n_{1},\ldots,n_{l}$. The following proposition provides an explicit expression for the product of unions of cycles.
@@
$\bigodot\limits_{i=1}^{l}C^{n_{i}}_{p_{i}}=C^{\frac{\tilde{p}_{l}\tilde{n}_{l}}{\lambda_{l}}}_{\lambda_{l}}\enspace.$
@@
We proceed by induction over $l$. First of all, we prove that the statement is true for $l=2$, i.e.,,
$C^{n_{1}}_{p_{1}}\cdot C^{n_{2}}_{p_{2}}=C^{\frac{(p_{1}p2)(n_{1}n_{2})}{\lambda_{2}}}_{\lambda_{2}}\enspace.$ (3)
Let us consider the case $n_{1}=n_{2}=1$. Since $C^{1}_{p_{1}}$ and $C^{1}_{p_{2}}$ can be viewed as finite cyclic groups of order $p_{1}$ and $p_{2}$, respectively, each element of the product of such cyclic groups has order $\lambda_{2}=\operatorname{lcm}(p_{1},p_{2})$ or, in other words, each element of $C^{1}_{p_{1}}\odot C^{1}_{p_{2}}$ belongs to some cycle of length $\lambda_{2}$. So, $C^{1}_{p_{1}}\odot C^{1}_{p_{2}}$ just consists of $(p_{1}\cdot p_{2})/\lambda_{2}$ cycles, all of length $\lambda_{2}$, and therefore
$C^{1}_{p_{1}}\cdot C^{1}_{p_{2}}=C^{\frac{p_{1}p_{2}}{\lambda_{2}}}_{\lambda_{2}}\enspace.$
@@
$C^{n_{1}}_{p_{1}}\cdot C^{n_{2}}_{p_{2}}=\sum\limits_{i=1}^{n_{1}}C^{1}_{p_{1}}\cdot\;\sum\limits_{j=1}^{n_{2}}C^{1}_{p_{2}}=\sum\limits_{i=1}^{n_{1}}\sum\limits_{j=1}^{n_{2}}(C^{1}_{p_{1}}\cdot\;C^{1}_{p_{2}})=\sum\limits_{i=1}^{n_{1}}\sum\limits_{j=1}^{n_{2}}C^{\frac{p_{1}p_{2}}{\lambda_{2}}}_{\lambda_{2}}=C^{\frac{(p_{1}p_{2})(n_{1}n_{2})}{\lambda_{2}}}_{\lambda_{2}}\enspace.$
Assume now that the equality holds for any $l>1$. Then, we get
$\sum\limits_{i=1}^{l+1}C^{n_{i}}_{p_{i}}=C^{\frac{\tilde{p}_{l}\tilde{n}_{l}}{\lambda_{l}}}_{\lambda_{l}}\cdot\;C^{n_{l+1}}_{p_{l+1}}=C^{\frac{\tilde{p}_{l}p_{l+1}\tilde{n}_{l}n_{l+1}}{\operatorname{lcm}(\lambda_{l},p_{l+1})}}_{\operatorname{lcm}(\lambda_{l},p_{l+1})}=C^{\frac{\tilde{p}_{l+1}\tilde{n}_{l+1}}{\lambda_{l+1}}}_{\lambda_{l+1}}\enspace,$
where $\lambda_{l+1}=\lambda(\lambda_{l},p_{l+1})$, $p_{l+1}=\tilde{p}_{l}p_{l+1}$ and $n_{l+1}=\tilde{n}_{l}n_{l+1}$. ∎
@@
$C^{1}_{p}\cdot X=C^{n}_{q}\enspace,$ (4)
where $X$ is the unknown and $p,n$ and $q$ are positive integers.
@@
   Question: Does Equation (4) admit any solution?
@@
Let $p,q,n$ be positive integers. Moreover, assume that the prime factorisations of $p,q$ and $n$ are:
@@
$q$ $=q_{1}^{k_{1}}\cdot\ldots\cdot q_{r}^{k_{s}}\text{ with }\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$
$n$ $=n_{1}^{m_{1}}\cdot\ldots\cdot n_{t}^{m_{t}}\text{ with }\operatorname{factors}\left(n\right)=\left\{n_{1},\ldots,n_{t}\right\}\enspace.$
The equation $C^{1}_{p}\cdot\sum_{i=1}^{l}C^{1}_{p^{\prime}_{i}}=C^{n}_{q}$ has solutions if and only if the two following conditions hold:
@@
    for all $i\in\left\{1,\ldots,r\right\}$ either $p_{i}^{h_{i}}\in\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$ or $p_{i}^{h_{i}}$ divides $n$.
@@
The left part of equation can be rewritten as $\sum_{i=1}^{l}C^{1}_{p}\cdot C^{1}_{p^{\prime}_{i}}$. Hence, by Proposition 1, $p$ must divide $q$. This means that $\operatorname{factors}\left(p\right)\subseteq\operatorname{factors}\left(q\right)$ and that for any $p_{i}^{h_{i}}$ there exists $j\in\left\{1,\ldots,s\right\}$ such that $q_{j}\in\operatorname{factors}\left(q\right)$ and $h_{i}\leq k_{j}$.
($\Leftarrow$) Given the factorisation of $q$ and Proposition 1, we can rewrite $C^{n}_{q}$ as follows.
$C^{n}_{q}=C^{1}_{q}\cdot C^{n}_{1}=C^{1}_{{q_{1}^{k_{1}}}\cdot\ldots\cdot{q_{s}^{k_{s}}}}\,\cdot C^{n}_{1}=C^{1}_{q_{1}^{k_{1}}}\cdot\ldots\cdot C^{1}_{q_{s}^{k_{s}}}\cdot C^{n}_{1}$
@@
$C^{1}_{q_{1}^{k_{1}}}\cdot\ldots\cdot C^{1}_{q_{s}^{k_{s}}}\cdot C^{n}_{1}=C^{1}_{q_{1}^{k_{1}}}\cdot\ldots\cdot C^{1}_{q_{s}^{k_{s}}}\cdot C^{1}_{{\mathsf{p}_{i}^{h_{i}}}}\cdot C^{\frac{n}{\left(p_{i}^{h_{i}}\right)}}_{1}.$
By repeating the same operation for all $p_{i}^{h_{i}}$ not included in $\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$, we find a solution since $\sum_{i=1}^{l}C^{1}_{p^{\prime}_{i}}$ will contain a cycle $C^{1}_{p^{\prime}_{i}}$ (with $p^{\prime}_{i}$ equals $q_{j}^{k_{j}}$) for any $p_{i}^{h_{i}}$ not included, and the set of self-loops.
($\Rightarrow$) Let us suppose that the equation $C^{1}_{p}\cdot\sum_{i=1}^{l}C^{1}_{p^{\prime}_{i}}=C^{n}_{q}$ holds for some $p^{\prime}_{1},p^{\prime}_{2},\ldots,p^{\prime}_{l}$. Let us assume that there exists $i$ such that $p_{i}^{h_{i}}\notin\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$. Then, $p_{i}^{h_{i}}$ will be equal to a certain $q_{j}^{y}$ with $y<k_{j}$. This implies that all $p^{\prime}_{1},p^{\prime}_{2},\ldots,p^{\prime}_{l}$ must contain exactly $q_{j}^{k_{j}}$ as a factor, since the $\operatorname{lcm}$ with $p$ must be $q$. Thus we can write:
$\sum_{i=1}^{l}C^{1}_{p^{\prime}_{i}}$ $=C^{1}_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C^{1}_{\frac{p^{\prime}_{i}}{\left(q_{j}^{k_{j}}\right)}}$
@@
$C^{1}_{p}\cdot\sum_{i=1}^{l}C^{1}_{p^{\prime}_{i}}$ $=C^{1}_{p}\cdot C^{1}_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C^{1}_{\frac{p^{\prime}_{i}}{\left(q_{j}^{k_{j}}\right)}}$
$=C^{1}_{\frac{p}{\left(q_{j}^{y}\right)}}\cdot C^{1}_{q_{j}^{y}}\cdot C^{1}_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C^{1}_{\frac{p^{\prime}_{i}}{\left(q_{j}^{k_{j}}\right)}}$
$=C^{1}_{\frac{p}{\left(q_{j}^{y}\right)}}\cdot C^{\left(q_{j}^{y}\right)}_{\left(q_{j}^{k_{j}}\right)}\cdot\sum_{i=1}^{l}C^{1}_{\frac{p^{\prime}_{i}}{\left(q_{j}^{k_{j}}\right)}}$
@@
Let $p,q$ be two positive integers. Let $p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}$ and $q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}$ be the prime number factorizations of $p$ and $q$, respectively. Let us also define $\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{r}\right\}$ and $\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$. Finally, let $\mathbf{F}$ be the set of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\setminus\operatorname{factors}\left(p\right)$. Algorithm 1 computes $\Pi_{\mathbf{F}}=\prod_{\mathbf{f}\in\mathbf{F}}\mathbf{f}$ without using the factorizations of $p$ and $q$. Moreover, the time complexity of Algorithm 1 is $O(\log p\cdot\log q)$.
@@
Denote $P=\operatorname{factors}\left(p\right)$ and $Q=\operatorname{factors}\left(q\right)$. If $g=\gcd(p,q)=1$ (Line $2$), then $P\cap Q=\emptyset$ and $\Pi_{\mathbf{F}}=q$.Hence, the algorithm returns $q$ (Line $4$). If $g\neq 1$, then taking $q/g$ keeps unchanged the $q_{j}^{k_{j}}$ belonging to $Q\setminus P$ and at the same time decreases of at least $1$ the exponents of the $q_{j}$ belonging to $P\cap Q$. Therefore, calling recursively (Line $6$) $\Pi_{\mathbf{F}}$ with $(q=q/g,g)$ keeps decreasing the exponents of $q_{j}$ belonging to $P\cap Q$ until they become 00. At that point, $\gcd(q,g)=1$ and the algorithm exits returning $\Pi_{\mathbf{F}}$. Remark that $\gcd(p,q)$ can be computed in $O(\log(\min(p,q)))$ and that the depth of the recursion is $O(\log q)$ in the worst case. Hence, the worst case complexity is $O(\log p\cdot\log q)$. ∎

### v4

@@ Abstract
Endowing the set of functional digraphs with the sum (disjoint union of digraphs) and product (standard direct product on digraphs) operations induces on FGs a structure of a commutative semiring $R$. The operations on $R$ can be naturally extended to the set of univariate polynomials $R[X]$ over $R$. This paper provides a polynomial time algorithm for deciding if equations of the type $AX=B$ have solutions when $A$ is just a single cycle and $B$ a set of cycles of identical size. We also prove a similar complexity result for some variants of the previous equation.
@@ 1 Introduction
A polynomial in $R[X_{1},X_{2},\ldots,X_{k}]$ represents a (infinite) set of functions which have a common substructure provided by the coefficients (modulo isomorphism). The factorization is an interesting inverse problem in this context. Indeed, assume to have a functional digraph $G$. Can $G$ be factorized i.e., decomposed into the product or the sum (or a combination of the two) of functional digraphs with smaller vertex set [^1] We stress that factorization here is meant in the context of abstract algebra which has nothing to do with the notion of factor graph often used in graph theory.? If this question has been largely investigated for general graphs, little is known for functional digraphs.
Let us consider another version of the factorization problem. Assume that we have partial information (or partial assumptions) on the factorization of $G$. This partial information is represented by a sequence of functional digraphs $A_{1},A_{2},\ldots,A_{k}$ and the question of factorization can be reformulated as “does the following equation
@@
admit a solution?”  And more generally one can ask for solutions of the following
@@
The interest of this result is manyfold. Without any doubt, it will significantly improve the performance of the software pipeline in  [4] for the case in which one is interested to the existence of solutions and not to their enumeration. It will shed some light in some connected important questions concerning the cancellation problem for functional digraphs i.e., the problem of establishing if $G\cdot F\approx G\cdot H$ implies $F\approx H$ where $F,G$ and $H$ are FGs and $\approx$ is the relation of graph isomorphism [5, 8]. We refer the reader to the classical book of Hammak for the factorization and the cancellation problems for more general graphs [7].
@@ 2 Background and basic facts
Figure 1: An example of polynomial equation over FGs (above) and a solution (below).
@@
It is therefore natural to ask if simpler equations admit solution algorithms with lower complexity. This will be very convenient in practical applications. Indeed, in Biology, when analyzing the behavior of a gene regulation network,for example, one can ask if the observed behavior is produced by the network at hand or if it is the result of the cooperation of simpler systems. These questions can be easily translated in terms of equations over FGs. However, gene regulation networks may contain hundreds of genes and hence effective methods are necessary to solve those equations.
@@ 3 Solving basic equations on permutations
From now on, we will focus on a subclass of functional digraphs, namely, the digraphs of permutations. For this reason, we introduce a convenient notation called C-notation inspired from [4]. We note $C_{p}$ the graph made by a single cycle of size $p$. Similarly, $n\cdot C_{p}$ (or simply $nC_{p}$) denotes a graph that is the sum of $n$ graphs $C_{p}$. According to this notation, it is clear that $aC_{p}+bC_{p}=(a+b)C_{p}$ and $a\cdot bC_{p}=abC_{p}$ for any natural $a,b$. With this notation, the classical result that the graph $G$ of a permutation is a union of disjoint cycles translates into
$G=\sum_{i=1}^{l}n_{i}C_{p_{i}}$
for suitable positive integers $p_{1},\ldots,p_{l}$ and $n_{1},\ldots,n_{l}$. Figure 2 provides an example of operations on permutation digraphs and their expression through the C-notation.
[…]
Figure 2: Example of operations on FGs of permutations. Using the C-notation, we have $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\cdot(\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}C_{3}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}C_{2})\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\cdot\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}C_{3}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\cdot\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=C_{6}+2C_{2}$.
The following proposition provides an explicit expression for the product of unions of cycles.
@@
$\prod\limits_{i=1}^{l}n_{i}C_{p_{i}}=\frac{\tilde{p}_{l}\tilde{n}_{l}}{\lambda_{l}}C_{\lambda_{l}}\enspace.$
@@
We proceed by induction over $l$. First of all, we prove that the statement is true for $l=2$, i.e.,
$n_{1}C_{p_{1}}\cdot n_{2}C_{p_{2}}=\frac{(p_{1}p_{2})(n_{1}n_{2})}{\lambda_{2}}C_{\lambda_{2}}\enspace.$ (3)
Consider the case $n_{1}=n_{2}=1$. Since $C_{p_{1}}$ and $C_{p_{2}}$ can be viewed as finite cyclic groups of order $p_{1}$ and $p_{2}$, respectively, each element of the product of such cyclic groups has order $\lambda_{2}=\operatorname{lcm}(p_{1},p_{2})$ or, in other words, each element of $C_{p_{1}}\cdot C_{p_{2}}$ belongs to some cycle of length $\lambda_{2}$. So, $C_{p_{1}}\cdot C_{p_{2}}$ consists only of $(p_{1}\cdot p_{2})/\lambda_{2}$ cycles, all of length $\lambda_{2}$, and therefore
$C_{p_{1}}\cdot C_{p_{2}}=\frac{p_{1}p_{2}}{\lambda_{2}}C_{\lambda_{2}}\enspace.$
@@
$n_{1}C_{p_{1}}\cdot n_{2}C_{p_{2}}=\sum\limits_{i=1}^{n_{1}}C_{p_{1}}\cdot\;\sum\limits_{j=1}^{n_{2}}C_{p_{2}}=\sum\limits_{i=1}^{n_{1}}\sum\limits_{j=1}^{n_{2}}(C_{p_{1}}\cdot\;C_{p_{2}})=\sum\limits_{i=1}^{n_{1}}\sum\limits_{j=1}^{n_{2}}\frac{p_{1}p_{2}}{\lambda_{2}}C_{\lambda_{2}}=\frac{(p_{1}p_{2})(n_{1}n_{2})}{\lambda_{2}}C_{\lambda_{2}}\enspace.$
Assume now that the equality holds for any $l>2$. Then, we get
$\sum\limits_{i=1}^{l+1}C^{n_{i}}_{p_{i}}=\frac{\tilde{p}_{l}\tilde{n}_{l}}{\lambda_{l}}C_{\lambda_{l}}\cdot\;n_{l+1}C_{p_{l+1}}=\frac{\tilde{p}_{l}p_{l+1}\tilde{n}_{l}n_{l+1}}{\operatorname{lcm}(\lambda_{l},p_{l+1})}C_{\operatorname{lcm}(\lambda_{l},p_{l+1})}=\frac{\tilde{p}_{l+1}\tilde{n}_{l+1}}{\lambda_{l+1}}C_{\lambda_{l+1}}\enspace,$
where $\lambda_{l+1}=\operatorname{lcm}(\lambda_{l},p_{l+1})$, $\tilde{p}_{l+1}=\tilde{p}_{l}p_{l+1}$ and $\tilde{n}_{l+1}=\tilde{n}_{l}n_{l+1}$. ∎
In the sequel, we will also make use of the following notation. Let $n$ be any integer strictly greater than $1$. We note $\operatorname{factors}\left(n\right)$ the set of prime factors of $n$ i.e., if $n=n_{1}^{m_{1}}\cdot\ldots\cdot n_{t}^{m_{t}}$, then $\operatorname{factors}\left(n\right)=\left\{n_{1},\ldots,n_{t}\right\}$ where $n_{1},\ldots,n_{t}$ are distinct primes.
@@
$C_{p}\cdot X=nC_{q}\enspace,$ (4)
where $X$ is the unknown and $p,n$ and $q$ are positive integers. We stress that the numbers $p,q$ and $n$ are just positive integers. They should not be thought of as prime integers unless explicitly mentioned. One of the issues of basic equations is that they might admit several distinct solutions. This is because some digraphs admit several distinct factorizations. Figure 3 provides an easy example of this fact.
[…]
Figure 3: Example of functional digraph having multiple factorizations. Using the C-notation, we have $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\cdot\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}C_{1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}C_{1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})\cdot\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}C_{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=2C_{2}$.
@@
   Question: Does the equation $C_{p}\cdot X=nC_{q}$ admit any solution?
@@
However, we are going to show that the problem can be solved in polynomial time in Theorem 4. The proof of Theorem 4 is based on the following characterization of the solutions of basic equations on the digraphs of permutations.
@@
Let $p,q,n$ be positive integers and let the prime decompositions of $p$ and $q$ be as follows:
@@
$q$ $=q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}\text{ with }\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$
The equation $C_{p}\cdot\sum_{i=1}^{l}C_{z_{i}}=nC_{q}$ has solutions (the values $z_{i}$ are the unknowns and need not to be distinct) if and only if the two following conditions hold:
@@
    $e$ divides $n$,
where $e=\prod_{v\in E}v$ with $E=\left\{q_{j}^{k_{j}}\;|\;\exists i\text{\ s.t.\@\ }p_{i}=q_{j}\text{ and }h_{i}<k_{j}\right\}$.
@@
The left part of equation can be rewritten as $\sum_{i=1}^{l}C_{p}\cdot C_{z_{i}}$. Hence, by Proposition 1, $p$ must divide $q$. This means that $\operatorname{factors}\left(p\right)\subseteq\operatorname{factors}\left(q\right)$ and that for any $p_{i}^{h_{i}}$ there exists $j\in\left\{1,\ldots,s\right\}$ such that $q_{j}\in\operatorname{factors}\left(q\right)$ and $h_{i}\leq k_{j}$.
($\Leftarrow$) Given the factorization of $q$ and Proposition 1, we can rewrite $nC_{q}$ as follows.
$nC_{q}=C_{q}\cdot nC_{1}=C_{{q_{1}^{k_{1}}}\cdot\ldots\cdot{q_{s}^{k_{s}}}}\,\cdot nC_{1}=C_{q_{1}^{k_{1}}}\cdot\ldots\cdot C_{q_{s}^{k_{s}}}\cdot nC_{1}$
@@
$C_{q_{1}^{k_{1}}}\cdot\ldots\cdot C_{q_{s}^{k_{s}}}\cdot nC_{1}=C_{q_{1}^{k_{1}}}\cdot\ldots\cdot C_{q_{s}^{k_{s}}}\cdot C_{{\mathsf{p}_{i}^{h_{i}}}}\cdot\frac{n}{\left(p_{i}^{h_{i}}\right)}C_{1}.$
Repeating the same operation for all $p_{i}^{h_{i}}$ not included in $\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$, we find a solution since $\sum_{i=1}^{l}C_{z_{i}}$ will contain a cycle $C_{z_{i}}$ (with $z_{i}$ equal to $q_{j}^{k_{j}}$) for any $p_{i}^{h_{i}}$ not included, and the set of self-loops.
($\Rightarrow$) Let us suppose that the equation $C_{p}\cdot\sum_{i=1}^{l}C_{z_{i}}=nC_{q}$ holds for some $z_{1},z_{2},\ldots,z_{l}$. Let us assume that there exists $i$ such that $p_{i}^{h_{i}}\notin\left\{q_{1}^{k_{1}},\ldots,q_{s}^{k_{s}}\right\}$. Then, $p_{i}^{h_{i}}$ will be equal to a certain $q_{j}^{y}$ with $y<k_{j}$. This implies that all $z_{1},z_{2},\ldots,z_{l}$ must contain exactly $q_{j}^{k_{j}}$ as a factor, since the $\operatorname{lcm}$ with $p$ must be $q$. Thus we can write:
$\sum_{i=1}^{l}C_{z_{i}}$ $=C_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C_{\frac{z_{i}}{\left(q_{j}^{k_{j}}\right)}}$
@@
$C_{p}\cdot\sum_{i=1}^{l}C_{z_{i}}$ $=C_{p}\cdot C_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C_{\frac{z_{i}}{\left(q_{j}^{k_{j}}\right)}}$
$=C_{\frac{p}{\left(q_{j}^{y}\right)}}\cdot C_{q_{j}^{y}}\cdot C_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C_{\frac{z_{i}}{\left(q_{j}^{k_{j}}\right)}}$
$=C_{\frac{p}{\left(q_{j}^{y}\right)}}\cdot q_{j}^{y}C_{q_{j}^{k_{j}}}\cdot\sum_{i=1}^{l}C_{\frac{z_{i}}{\left(q_{j}^{k_{j}}\right)}}$
@@
Theorem 3 provides a condition to decide whether a solution to a basic equation can exist. However, we are not going to exploit it directly, since it requires the prime factorization of the integers $p$ and $q$ which would imply unnecessarily high complexity. In fact, the algorithms we are going to conceive will essentially use simple arithmetical operations or basic functions on the integers. In this regard, we consider that the worst-case time complexity of the addition (resp., multiplication) between two integers is $O(n)$ (resp., $O(n^{2})$), where $n$ is the input size in bits. The division between two integers is assumed to have the same complexity as the multiplication. Concerning $\gcd$, we consider the classical Euclidean GCD algorithm which has worst-case time complexity $O(n^{2})$.
@@
Let $p,q$ be two positive integers. Let $p_{1}^{h_{1}}\cdot\ldots\cdot p_{r}^{h_{r}}$ and $q_{1}^{k_{1}}\cdot\ldots\cdot q_{s}^{k_{s}}$ be the prime number factorizations of $p$ and $q$, respectively. Let us also define $\operatorname{factors}\left(p\right)=\left\{p_{1},\ldots,p_{r}\right\}$ and $\operatorname{factors}\left(q\right)=\left\{q_{1},\ldots,q_{s}\right\}$. Finally, let $\mathbf{F}$ be the set of all $q_{j}^{k_{j}}$ such that $q_{j}\in\operatorname{factors}\left(q\right)\setminus\operatorname{factors}\left(p\right)$. Algorithm 1 computes $\Pi_{\mathbf{F}}=\prod_{\mathbf{f}\in\mathbf{F}}\mathbf{f}$ without using the factorizations of $p$ and $q$. Moreover, the worst-case time complexity of Algorithm 1 is $O(n^{3})$, where $n$ is the size of the input in bits.
@@
Denote $P=\operatorname{factors}\left(p\right)$ and $Q=\operatorname{factors}\left(q\right)$. If $g=\gcd(p,q)=1$ (Line $2$), then $P\cap Q=\emptyset$ and $\Pi_{\mathbf{F}}=q$. Hence, the algorithm returns $q$ (Line $4$). If $g\neq 1$, then taking $q/g$ keeps unchanged the $q_{j}^{k_{j}}$ belonging to $Q\setminus P$ and at the same time decreases of at least $1$ the exponents of the $q_{j}$ belonging to $P\cap Q$. Therefore, calling recursively (Line $6$) $\Pi_{\mathbf{F}}$ with $(q=q/g,g)$ keeps decreasing the exponents of $q_{j}$ belonging to $P\cap Q$ until they become 00. At that point, $\gcd(q,g)=1$ and the algorithm exits returning $\Pi_{\mathbf{F}}$. Remark that at each call of $\Pi_{F}$, we compute a $\gcd$ (Line $2$) plus a division (Line $6$) which cost $O(n^{2})$ in total. Since the depth of the recursion is $O(n)$, the worst-case time complexity for the algorithm is $O(n^{3})$. ∎